OpenStudy (anonymous):

Find the quotient and the remainder of 3x^3 - 2x^2 + 4x-3 / x + 4

OpenStudy (cwrw238):

you can use long division or synthetic division to find the quotient and remainder also you can find the remainder using the remainder theorem as follows plug in x = -4 ( as you are dividing by (x+4): 3(-4)^3 -2(-4)^2 +4(-4) - 3 = -192 + 32 - 16 - 3 = -179

OpenStudy (cwrw238):

i made a mistake using remainder theorem : its -192 - 32 - 16 - 3 = -243

OpenStudy (cwrw238):

and another mistake in long division - sorry guy - not my best day today!

OpenStudy (cwrw238):

have you done long division or synthetic division already?

OpenStudy (cwrw238):

http://www.purplemath.com/modules/synthdiv.html - this explains it pretty well although i always use long division with polynomials because i'm used to it

OpenStudy (cwrw238):

hope this is of help to you